OpenStudy (anonymous):

Determine which system below will produce infinitely many solutions. -4x + 6y = 20 4x - 3y = -2 -2x + 4y = 18 4x - 8y = -36 3x - 2y = 5 -6x - 4y = -10 5x - 3y = 12 15x - 9y = 24

OpenStudy (igreen):

Infinitely many solutions = Same line

OpenStudy (twopointinfinity):

I'm gonna determine how you should put a shirt on.

OpenStudy (cwrw238):

if one of the pair of equations can be reduced to the other there are infinite solution

OpenStudy (anonymous):

Lol @TwoPointInfinity

OpenStudy (anonymous):

I think it's cute :)

OpenStudy (hauntedwoodsgal):

lol

OpenStudy (anonymous):

who

OpenStudy (anonymous):

help

OpenStudy (anonymous):

I'm not sure if I'm right but would it be 5x - 3y = 12

OpenStudy (cwrw238):

example x + y = 2 2x + 2y = 4 - the second equation can be formed by multiplying the first by 2 plotting them will give the same line

OpenStudy (anonymous):

A -4x + 6y = 20 4x - 3y = -2 B -2x + 4y = 18 4x - 8y = -36 C 3x - 2y = 5 -6x - 4y = -10 D 5x - 3y = 12 15x - 9y = 24

OpenStudy (anonymous):

islike that

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

are u sure

OpenStudy (cwrw238):

check out B

OpenStudy (anonymous):

So everything is good but B

OpenStudy (cwrw238):

write them in different way 8y = 4x + 36 4y = 2x + 18

OpenStudy (anonymous):

so is b

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome! :)